### v1

# Feedback Descent: Open-Ended Text Optimization via Pairwise Comparison

Abstract
We introduce Feedback Descent, a framework that optimizes text artifacts—prompts, code, and molecules—through structured textual feedback rather than relying solely on scalar rewards. By preserving detailed critiques instead of compressing them to binary preferences, Feedback Descent widens the information bottleneck in preference learning, enabling directed optimization in text space rather than weight space. We show that in-context learning can transform structured feedback into gradient-like directional information, enabling targeted edits. Unlike prior approaches that collapse judgments into single bits, our evaluators pair each comparison with textual feedback, which functions as high-bandwidth supervision. The iteration loop is done purely at inference time, without modifying any model weights, and is task-agnostic. We evaluate Feedback Descent on three diverse domains and find that it outperforms state-of-the-art prompt optimization (GEPA), reinforcement learning methods (GRPO, REINVENT), and even specialized graph-based molecular optimizers. In the DOCKSTRING molecule discovery benchmark, Feedback Descent identifies novel drug-like molecules surpassing the $99.9$th percentile of a database with more than 260,000260{,}000 compounds across six protein targets.
1 Introduction
A central goal of machine learning is building systems that can perform tasks beyond human capabilities. Reinforcement learning is a powerful framework that accomplishes this goal, since it can optimize with respect to feedback on its own outputs, rather than relying on supervised examples of desired outputs. Indeed, recent language models have demonstrated impressive feats in domains like math and programming [openai2024o1, deepseek2025r1, alphageometry2025, zhu2024deepseek] through a combination of reinforcement learning and text-based reasoning. Unfortunately, existing reinforcement learning frameworks are designed to learn from impoverished supervision signals, typically either scalar rewards or pairwise preference data, where each annotation conveys at most a single bit per pair. These bottlenecks discard information about why one behavior is better and how to improve—information available in environment feedback or easily elicited from humans during annotation [wu2023finegrained, just2024datacentric].
Our goal is to widen this information bottleneck, i.e., significantly increase the information the system can extract per unit of experience [silver2025era]. Collecting more detailed feedback is straightforward, e.g., with brief rationales explaining preferences; the challenge is turning such feedback into measurable improvement. Because free-form feedback does not define a differentiable objective, it cannot directly drive weight updates via backpropagation. Our approach iterates at inference time, using language models to translate accumulated feedback into targeted edits of text artifacts (prompts, code, molecules, JSON configs) that improve a final performance objective, without any weight updates.
To that end, we introduce Feedback Descent, a framework for continual optimization in text space. At each iteration, we prompt a language model to propose an improved version of the current best artifact, conditioned on all previous feedback. We compare this candidate against the current best, and the evaluator returns a preference along with textual feedback explaining the choice. If the candidate is preferred, it becomes the new best. Repeating this loop yields semantically local, feedback-aligned improvements that implement gradient-like steps in text space. See Figure˜1 for a conceptual illustration. We provide theoretical intuition for why Feedback Descent can be effective. Under appropriate assumptions about feedback quality and problem structure, we demonstrate that textual feedback can provide directional information, enabling efficient optimization.
Our contributions are threefold. First, we formalize why textual feedback enables dimension-free convergence while zeroth-order methods suffer exponential slowdown with effective dimensionality, identifying when and why structured feedback outperforms scalar rewards. Second, we demonstrate cross-domain generality: Feedback Descent works across three qualitatively distinct domains (visual design, prompt optimization, molecule design) with the same iterative loop. Third, we validate competitive or superior performance versus specialized methods, achieving competitive results with the state-of-the-art in prompt optimization (GEPA) while outperforming a reinforcement learning baseline (GRPO). In the molecule design experiment, Feedback Descent outperforms specialized molecular optimizers (Graph MCTS/GA, REINVENT) despite operating purely on text representations, discovering molecules that surpass the 99.9th percentile of a 260,000-compound database of molecules.
Figure 1: A conceptual illustration of feedback descent. At each iteration, we compare the previous best artifact with a new candidate. The evaluator provides both a pairwise preference and textual feedback. Preferences ensure the selection of better candidates, while feedback accumulates directional information that guides semantically meaningful edits.
2 Feedback Descent: Open-Ended Text Optimization
We propose Feedback Descent, a framework for open-ended optimization of text-representable artifacts whose quality is easier to judge than to construct. Feedback Descent converts comparative textual feedback into directed semantic edits and iterates in a self-improvement loop. As a running example, consider optimizing SVG code to render better images of a unicorn. Current vision-language models can reliably compare two renderings and explain the choice, even if writing high-quality SVG from scratch is difficult. Through Feedback Descent, we can convert these explanations into directed edits that aim to produce an artifact that surpasses all previous ones.
2.1 Problem Setup
Let $\mathcal{S}$ be the space of token sequences, and let $x\in\mathcal{S}$ denote an artifact (e.g., SVG code). Given the current best $x_{t}^{\star}\in\mathcal{S}$ and a candidate $x\in\mathcal{S}$, the evaluator returns
$\mathsf{E}(x,x_{t}^{\star})\to\big(p\in\{0,1\},\ r\in\mathcal{S}\big),$ (1)
where $p=1$ indicates $x\succ x_{t}^{\star}$ and $r$ is a textual feedback explaining why the winner is better and how to improve. We append $r_{t}$ to a history $\mathcal{R}_{t}=\{(x_{1},r_{1}),\ldots,(x_{t},r_{t})\}$ and iterate, keeping track of the current best artifact $x_{t}^{\star}$.
2.2 Feedback Descent
Feedback Descent operates as an iterative optimization loop that maintains a single best artifact $x_{t}^{*}$ and progressively improves it through feedback-guided mutations and comparative evaluation. Throughout, we use $\mathcal{M}$ to denote the language model used for generating improved candidates.
Initialization and termination. We initialize $x_{0}^{*}$ by prompting a language model with the task description alone (e.g., "Generate SVG code for a unicorn"). The algorithm runs for a fixed budget of $T$ iterations or until convergence (defined as no improvement for $k$ consecutive iterations).
Proposing semantic mutations via prompting. The mutation step leverages a language model’s in-context learning capabilities. At iteration $t$, we prompt the model with the current best artifact $x_{t}^{*}$ and accumulated feedback $\mathcal{R}_{t-1}$ to generate an improved candidate:
$x_{t}=\mathcal{M}\left(x_{t}^{*},\mathcal{R}_{t-1}\right)$ (2)
The prompt instructs the model to address previous critiques while preserving successful elements. These prompts are intentionally minimal: the optimization signal comes from the accumulated feedback rather than heavy prompt engineering.
Algorithm 1 Feedback Descent
1:Initial text $x_{0}$, Language model $\mathcal{M}$, $T$
[…]
3:for $t=1$ to $T$ do
4:  $x_{t}\leftarrow\mathcal{M}(x^{*},\mathcal{R})$ $\triangleright$ Propose (2)
5:  $p_{t},r_{t}\leftarrow\textsc{Compare}(x_{t},x^{*})$ $\triangleright$ Compare (1)
6:  $\mathcal{R}\leftarrow\mathcal{R}\cup\{(x_{t},r_{t})\}$
7:  if $p_{t}=1$ then
8:   $x^{*}\leftarrow x_{t},\mathcal{R}\leftarrow\emptyset$ $\triangleright$ Update + reset
9:return $x^{*}$
They include basic task context, the current best artifact, and feedback from previous comparisons. Complete prompt templates for each domain are provided in Section˜C.3.
Selection and update. We compare the new candidate $x_{t}$ against the current best $x_{t}^{*}$ using the evaluator $\mathsf{E}(x_{t},x_{t}^{*})$, which returns both a binary preference $p_{t}$ and a textual feedback $r_{t}$. In our running SVG example, examples of feedback include “adjust the stroke width”, “make sure the legs are connected to the body”, and “add a shadow to the unicorn’s mane”. Regardless of the preference outcome, we always add the feedback to our history: $\mathcal{R}_{t+1}=\mathcal{R}_{t}\cup\{(x_{t},r_{t})\}$. If $p_{t}=1$ (candidate is preferred), we update $x_{t+1}^{*}=x_{t}$; otherwise we keep $x_{t+1}^{*}=x_{t}^{*}$. We summarize the overall process in Algorithm˜1.
2.3 Analogy to Gradient Descent
The key algorithmic insight is best understood by analogy to the gradient descent algorithm. Just as gradients provide the direction of steepest ascent under local linearity, textual feedback can suggest plausible directions of improvement in semantic space. For our SVG example, if the feedback indicates “needs more defined horn shape,” we expect that a small edit to the horn shape that preserves overall structure will likely be an improvement.
Of course, textual feedback is not a literal gradient. It is approximate and occasionally contradictory—optimization with such feedback does not have convergence guarantees in the same way that gradient descent does. Instead, feedback acts as a heuristic directional cue, offering higher-bandwidth supervision than a binary preference signal or a scalar reward, just as first-order optimization is fundamentally faster than zeroth-order optimization [nemirovski1983problem, agarwal2011information, nesterov2017random]. We hypothesize that an open-ended optimization loop based on such cues can succeed, supported by prior evidence that language models reliably translate textual instructions into concrete modifications. Examples include generating code changes [chen2021evaluating, austin2021program, nijkamp2022codegen, wang2023codet5+, roziere2023code, guo2024deepseek, lozhkov2024starcoder2, codegemma2024], following complex multi-step instructions [ouyang2022training, wei2022finetuned, chung2022scaling, longpre2023flan, zhang2024instruction], targeted text modifications [schick2022peer, du2022read, madaan2023selfrefine, welleck2023generating, mishra2023help], and decomposing high-level goals into executable action sequences [schick2023toolformer, parisi2022talm, yao2022react, qin2023tool, wang2024llm, agarwal2025toolrm].
We emphasize that this gradient analogy serves only as an intuition: gradients indicate directions in continuous parameter space, while rationales indicate directions in semantic space, without guarantees of smoothness or convergence. Our experiments in Section˜5 demonstrate that feedback-guided optimization indeed produces consistent improvements across tasks, validating that such heuristic directional cues are sufficient to drive open-ended text optimization.
3 Directional Information is Critical When Effective Dimensionality is High
In this section, we give intuition for why and when Feedback Descent is particularly effective, using a minimal idealized model of directional information based on comparisons with feedback. We provide simplified statements here, and show full proofs in Appendix˜A.
Pairwise preference alone is a binary signal: it tells us which artifact is better but not how to improve. Adding a short rationale (“legs disconnected from body; increase stroke width”) turns each comparison into a directional cue in semantic space. Iterating this loop accumulates such cues, allowing a mutator LLM to make small, targeted edits that are more likely than chance to increase quality. Even when individual rationales are imperfect, the aggregate message across failures (what kept losing and why) continually refines the direction of improvement.
As an idealized model, map each text $x$ to a latent feature $z=\phi(x)\in\mathbb{R}^{d}$ with utility $r(z)=R(x)$, where $d$ denotes the effective dimensionality—the number of independent directions along which artifacts can vary meaningfully. Let $z^{\star}=\arg\max_{z}r(z)$ be the optimum. For problems of interest, $d$ may be in the hundreds or thousands, reflecting the many degrees of freedom in text, molecule, or image design spaces.
A rationale $r_{t}$ is useful if its induced edit defines a direction $v_{t}$ positively correlated with the gradient $\nabla r(z_{t})$. Under this model, feedback carries gradient information and systematically reduces error, yielding linear convergence in expectation, even with imperfect rationales.
Proposition 1 (Directional updates achieve dimension-free convergence, informal).
Suppose we perform $N$ rationale-guided updates, where each update direction $v_{t}$ is on average positively aligned with $\nabla r(z_{t})$. Then the expected error after $N$ evaluations satisfies
$\mathbb{E}[\,r(z^{\star})-r(z_{N})\,]\;\leq\;(1-\alpha)^{N}\,[r(z^{\star})-r(z_{0})],$
for some constant $\alpha>0$ depending only on the alignment and variability of the feedback. Crucially, this rate does not depend on the dimension $d$.
While rationale-guided updates exploit directional information to achieve convergence rates that do not depend on the ambient dimension, methods that rely only on function evaluations or binary preferences suffer severe dimension-dependent slowdowns:
Proposition 2 (Inefficiency of zeroth-order search, informal).
Consider zeroth-order methods with $N$ evaluations in $\mathbb{R}^{d}$. For a strongly concave quadratic, to reach $\epsilon$-accuracy one needs at least $N\;\geq\;(c/\epsilon)^{d/2}$ and best-of-$N$ random sampling achieves at best
$\mathbb{E}[\,r(z^{\star})-r(z_{N})\,]=\Omega(N^{-2/d}).$
Thus, while rationale-guided updates drive $r(z_{N})$ toward $r(z^{\star})$ at a dimension-free linear rate, zeroth-order search suffers exponential slowdowns as $d$ grows.
We emphasize that these propositions are highly idealized, serving only to illustrate a general principle: in high-dimensional search, even weak directional cues can accumulate into steady improvement, whereas unguided search scales exponentially worse. Thinking in terms of a continuous latent space is a pragmatic abstraction; modern ML operates on continuous weights and embeddings that approximate local structure even in discrete domains. By the same logic, many forms of human progress, including writing, coding, and science, often arise from successive refinements within an implicit semantic space shared by experts. The following statements are for motivational clarity, not literal models of the systems in Section˜5.
The propositions rest on well-known results in optimization and combinatorics. Our contribution is applying this intuition to open-ended text optimization, where it highlights effective dimensionality as the key bottleneck. To our best knowledge, the closest prior work that adopts this framing is the empirical taxonomy and optimizer design of nie2024importance; building on their study, we identify effective dimensionality as the key limiting factor, showing that in open-ended settings where artifacts admit many possible alterations, even weak but aligned textual rationales provide the directional information needed for linear convergence, while zeroth-order methods degrade exponentially.
Implications. First, even imperfect cues accelerate optimization when aggregated, since their consistent positive correlation with the gradient accumulates into reliable directional information. Second, in problems with high effective dimensionality, where exploration can proceed in unbounded directions, pure zeroth-order search becomes extremely inefficient. In this regime, obtaining directed feedback on the most salient aspects, whether from experimental results, an expert’s feedback, or exploiting a generator-verifier gap, is essential for sustained progress. This is especially important for open-ended problems where the optimum is not known in advance.
Directional Information is Advantageous in High Dimensions Unlike zeroth-order methods, which scale poorly with dimension, Feedback Descent remains effective when the effective dimensionality is high because feedback provides directional cues that guide optimization.
4 Related Work
Preference Learning. Preference learning methods learn from pairwise comparisons [christiano2017deep, ouyang2022training, azar2023general, ethayarajh2024kto, munos2024nash]; recent advances include bypassing the need for a reward model [rafailov2023direct], iterative optimization under KL constraints [xiong2023iterative], and adaptive scaling techniques [wang2024adaptive]. However, these methods fundamentally compress complex human reasoning into binary or scalar preferences, thereby forgoing the rich explanatory content that humans can naturally provide alongside their judgments [wirth2017survey]. Recent work shows that fine-grained feedback significantly improves reward modeling [wu2023finegrained, yu2024selfgenerated], and incorporating rationales alongside preferences provides richer training signals [just2024datacentric]. Unlike prior work that uses rationales to improve weight-based training, we leverage detailed textual feedback entirely at inference time, widening the information bottleneck without requiring retraining.
Gradient-Free Optimization. Zeroth-order optimization methods [chen2019zoadamm] operate without access to gradients, typically achieving convergence rates that scale poorly with dimension. While the black-box nature of modern LLMs has spurred interest in applying gradient-free approaches [han2023match, sun2024bbtv2, jin2024derivative, brahmachary2024evolution], these methods face fundamental challenges in high-dimensional spaces. Our approach explores whether textual rationales can provide useful directional information for optimization, similar to how nie2024importance shows that LLMs can be effective optimizers when provided with directional feedback from historical traces.
Optimizing Compound AI Systems. Compound AI systems, i.e., modular architectures involving multiple LLM invocations and complex control flow, such as agents or scaffolding techniques [yao2022react], present unique optimization challenges due to their modularity. Several approaches have emerged to tackle this complexity, including optimization for searching and bootstrapping few-shot in-context examples [khattab2022dspy, khattab2024dspy, opsahl2024optimizing] and reflective prompt evolution [agrawal2025gepa]. The closest prior work to Feedback Descent is TextGrad [yuksekgonul2024textgrad], which proposes PyTorch-like framework for automatic differentiation via text, backpropagating textual gradients through computation graphs. A core difference is that TextGrad optimizes “pointwise”, i.e., it proposes a new instance of the artifact based only on the latest one. In contrast, Feedback Descent keeps an explicit trajectory-level buffer of comparative feedback. As we will see in Section˜5, Feedback Descent is much more scalable at long-horizon optimization compared to TextGrad.
Figure 2: Iterative progression of SVG unicorn optimization under the realism judge. Feedback Descent produces gradual, semantically meaningful improvements through accumulating directional cues.
Inference-Time Optimization for LLMs. Inference-time optimization improves performance without weight updates by performing additional computation at generation. This paradigm includes self-critique and refinement cycles (constitution-guided critique [bai2022constitutional]; Self-Refine [madaan2023selfrefine]) test-time scaling via best-of-$N$, multi-step reasoning, and tree search [cobbe2021training, zelikman2022star, yao2023tree], and iterative prompt optimization [zhou2022ape, yang2023opro, pryzant2023protegi]. Several works report that strategically allocating inference-time compute yields large gains [snell2024scaling, muennighoff2025s1, geiping2025scaling, zhou2025thinking].
We build on the growing consensus that natural language is a particularly powerful medium for inference-time improvement. Natural language traces enable models to reason effectively in complex environments [lampinen2022tell, wei2022chain], and language models can reliably map textual instructions to concrete modifications [chen2021evaluating, austin2021program, saunders2022self, scheurer2023training]. However, existing methods often rely on random sampling of self-generated critiques, which may be noisy or fail to capture external preferences. In contrast, we leverage external rationales as directional information, enabling guided search in the semantic space.
5 Experiments
We evaluate Feedback Descent across three diverse domains—visual design, prompt optimization, and molecule discovery—to demonstrate its generality and effectiveness. Our evaluation spans diverse representations and evaluation modalities: (1) Representation diversity: SVG (spatial/geometric), prompts (instructional text), molecules (chemical structures). (2) Evaluation modality: SVG uses vision-language model comparison, prompt optimization uses dataset-specific accuracy metrics, molecules use computational docking scores. Together, these answer: Can a single optimization framework, with no domain-specific engineering, match or exceed specialized methods purely through structured feedback?
5.1 Experimental Domains
We describe each evaluation domain and how we obtain pairwise comparisons augmented with textual rationales.
Figure 3: Example images generated by Feedback Descent under six different judge criteria. Feedback Descent yields visually distinct objects aligned with the aesthetic criteria preferred by each judge.
Setup Subject Ink Wash Minimalist Realism Retro Arcade Stained Glass
From Scratch City Skyline 100.0% 100.0% 100.0% 100.0% 100.0%
Mushroom 100.0% 100.0% 83.3% 100.0% 100.0%
Robot 100.0% 100.0% 100.0% 100.0% 100.0%
Unicorn 100.0% 100.0% 100.0% 100.0% 100.0%
Judge-Aware City Skyline 50.0% 100.0% 100.0% 100.0% 88.9%
Mushroom 100.0% 100.0% 100.0% 100.0% 94.1%
Robot 100.0% 100.0% 100.0% 93.8% 100.0%
Unicorn 89.5% 100.0% 100.0% 93.8% 80.0%
Table 1: Win rates after five iterations comparing Feedback Descent against direct prompting under two conditions: from Scratch and Informed of the judge rubric. Values above 50% indicate that Feedback Descent outperforms the baseline; denotes the only failure case where direct prompting wins. Iterative feedback consistently improves SVG designs over direct prompting.
SVG optimization. Taking inspiration from bubeck2023sparks, we evaluate the ability of models to output SVG code for illustrations of unicorns. We use a set of five diverse judge prompts, each preferring a different aesthetic: ink wash painting style, minimalist design, realism, retro arcade pixel-art motifs, and stained glass artwork. We compare rendered SVGs using GPT-5-mini, which outputs both a binary preference and short textual feedback. To mitigate order bias, we perform two judgments with swapped image orders (A-B and B-A) and declare a winner only if both judgments are consistent. Otherwise, we try again, up to three times, and discard if no consistent winner emerges.
Prompt optimization. We follow the setup of GEPA [agrawal2025gepa] across four diverse tasks: multi-hop reasoning (HotpotQA; yang2018hotpotqa), instruction following (IFBench; pyatkin2025generalizing), privacy-aware delegation (PUPA; li2025papillon), and retrieval-augmented verification (HoVer; jiang2020hover). We evaluate on both open-source (Qwen3-8B; yang2025qwen3) and proprietary (GPT-4.1 mini) models. For each task, we use the same multi-stage programs from GEPA, where the number of stages differs across datasets, and we jointly optimize the prompts for all stages using Feedback Descent. Optimization is driven by training examples: candidate prompts are updated based on performance on the training set and textual feedback describing which constraints were satisfied or violated. All candidate prompts are scored on validation examples, and the prompt with the highest validation accuracy rate is selected. We report performance on held-out test examples.
Molecule discovery. We evaluate on molecular docking tasks using DOCKSTRING [garciaortegon2021dockstring] docking scores and drug-likeness (QED). DOCKSTRING provides a realistic drug discovery setting where molecules are evaluated based on their predicted binding affinity to medically relevant targets, rather than relying solely on simple physicochemical properties. We focus on challenging optimization tasks across six protein targets: ADRB1, PGR, PPARA, PPARG, CDK2, and F2. Following DOCKSTRING, we compute the combined score $s=-\textrm{Vina}-10\times(1-\textrm{QED})$. We represent molecules as SMILES strings [weininger1988smiles] and evaluate using DOCKSTRING’s molecular docking pipeline to compute Vina scores (binding affinity). The feedback system provides rich structured information, including RDKit molecular descriptors [landrum2006rdkit], similarity searches against known compounds from molecular databases [liu2007bindingdb, gilson2016bindingdb, gaulton2012chembl, mendez2019chembl], and detailed docking results. In the system prompt, we also provide the LLM information about the protein target obtained from the UniProt database [uniprot2023]. Together, this provides the LLM with detailed feedback on molecular properties that affect binding affinity, drug-likeness violations, and comparisons to known active compounds.
5.2 SVG Optimization
We evaluate iterative feedback against direct prompting across two generators, GPT-4o-mini and GPT-5-mini. The direct prompting baseline receives the full evaluation rubric and is tasked with producing a single best design. Feedback Descent instead begins with an initial set of candidates, and through 5 rounds of structured feedback and improvement, refines designs using judge comparisons that reflect aesthetic criteria. We test two initialization regimes: Scratch, which starts from images simply instructed to generate images of unicorns, and Informed, which starts from the strongest direct generations conditioned on the rubric, determined by the LLM judge.
Method Qwen3-8B GPT-4.1 Mini
HpQA IFBench Hover PUPA HpQA IFBench Hover PUPA
DSPy Default [khattab2024dspy] 42.33 36.90 35.33 80.82 38.00 47.79 46.33 78.57
MIPROv2 [opsahl2024optimizing] 55.33 36.22 47.33 81.55 58.00 49.15 48.33 83.37
GRPO [shao2024deepseekmath] 43.33 35.88 38.67 86.66
GEPA [agrawal2025gepa] 62.33 38.61 52.33 91.85 69.00 52.72 51.67 94.47
GEPA+Merge [agrawal2025gepa] 64.33 28.23 51.67 86.26 65.67 55.95 56.67 96.46
Feedback Descent (ours) 60.00 38.78 60.00 90.90 68.33 54.59 57.67 85.66
Table 2: Prompt optimization results across multiple benchmarks. GEPA consistently outperforms baseline methods across different model architectures.
Results. Table˜1 shows the win rates after 5 iterations. For both GPT-4o-mini and GPT-5-mini, Feedback Descent reliably improves outputs over the initial population. Furthermore, qualitative examples in Figure˜3 demonstrate that the procedure consistently produces unicorns whose visual style diverges across judges, aligning with aesthetic criteria such as geometry, minimalism, or retro arcade motifs.
Iterative feedback can elicit better outputs from the same model Because of a generator-verifier gap, even prompting with the exact judge rubric is suboptimal for SVG generation. Feedback Descent elicits better images from the same generator by iteratively proposing improvements guided by feedback.
5.3 Prompt Optimization
We compare Feedback Descent against five baselines: the default prompt implemented in the DSPy program [khattab2024dspy, Default], a Bayesian optimization approach for selecting instructions and demonstrations [opsahl2024optimizing, MIPROv2], online reinforcement learning [shao2024deepseekmath, GRPO], and a reflective prompt evolution method [agrawal2025gepa, GEPA]. All baselines are run under matched rollout budgets for fair comparison, and the reported baseline results are from agrawal2025gepa.
Each example produces pointwise feedback about which constraints were satisfied or violated. To construct the pairwise feedback for Feedback Descent, we stratify the examples into quadrants based on whether each prompt resulted in a correct response. We then ask the model to propose textual descriptions of inputs where these discrepancies arise. We then statistically validate each hypothesis, filtering for ones that correspond to consistent differences in performance between the prompts. This process distills the true global differences between the two prompts.
Table˜2 shows that Feedback Descent is competitive with GEPA across both models, achieving the best performance on IFBench and Hover, while GEPA leads on HotpotQA and PUPA. Despite GEPA being specifically engineered for prompt optimization with coordinate descent and Pareto frontier maintenance, Feedback Descent achieves competitive performance with a simpler approach: jointly optimizing all prompts at once via automated textual summaries of pairwise performance differences. We do not claim to present a state-of-the-art prompt optimizer; rather, these results demonstrate that our general-purpose framework remains competitive with specialized methods while requiring minimal domain-specific engineering.
Pairwise Comparisons Suffice for Competitive Prompt Optimization Feedback Descent remains competitive with GEPA, a state-of-the-art prompt optimization method, despite employing a simpler, domain-agnostic approach that focuses on summarizing pairwise performance differences.
Method ADRB1 PGR PPARA PPARG CDK2 F2
DOCKSTRING (N=260155)
Top $50\%$ 5.305 3.478 4.549 4.210 4.385 4.168
Top $90\%$ 8.785 7.878 7.987 7.658 7.733 7.477
Top $99\%$ 9.620 8.703 8.718 8.449 8.453 8.139
Top $99.9\%$ 10.209 9.260 9.230 9.012 8.979 8.722
Top $99.99\%$ 10.742 9.723 9.821 9.518 9.509 9.252
Best Molecule 11.330 9.742 9.907 9.529 9.534 9.311
GP-BO [tripp2021a] 10.552 9.307 9.680 9.485 9.067 8.686
Graph MCTS [jensen2019graph] 8.883 7.819 7.363 7.134 7.777 6.310
Graph GA [jensen2019graph] 9.145 8.670 8.598 8.327 8.288 8.102
SMILES GA [brown2019guacamol] 9.334 8.335 9.052 8.560 8.268 7.984
REINVENT [olivecrona2017molecular] 9.018 8.267 8.430 8.347 8.226 8.139
TextGrad [yuksekgonul2024textgrad] 8.531 8.057 7.953 7.256 8.174 7.357
Feedback Descent 10.623 9.615 9.919 10.187 9.803 9.300
Table 3: Comparison of molecule optimization methods on six protein targets. Fragment-based algorithms (denoted by $\dagger$) operate directly on molecular graphs, giving them structural priors unavailable to purely text-based methods. For each target, the top generative result is in bold, and any population in the DOCKSTRING that exceeds the best generative result is underlined. Feedback Descent rivals or surpasses specialized molecular optimizers across all six targets.
Figure 4: Scatter plots of Tversky similarity to approved drugs against docking scores, showing weak or negative correlations across targets. High-scoring molecules discovered by Feedback Descent are far from any known drugs.
5.4 Molecule Optimization (DOCKSTRING)
We compare against baselines implemented in the mol_opt repository [gao2022molopt], Our comparisons include a genetic algorithm [brown2019guacamol, SMILES GA], reinforcement learning [olivecrona2017molecular, REINVENT], fragment-based algorithms [jensen2019graph, Graph MCTS/GA], and Bayesian optimization on molecular graphs [tripp2021a, GP-BO]. Because fragment-based methods exploit graph-level structural priors, the most direct comparison is to the text-only baselines: SMILES-GA and REINVENT. We also compare our approach with TextGrad [yuksekgonul2024textgrad], a recent work that similarly utilizes an LLM to make textual updates to SMILES strings. The key difference is that TextGrad’s improvement proposals are pointwise, conditioning only on the latest molecule, whereas Feedback Descent conditions on accumulated feedback history, enabling more effective continual improvement at high iteration budgets.
Results. Table˜3 summarizes optimization outcomes across six protein targets. Feedback Descent outperforms all baselines and achieves the strongest scores on all targets. On multiple proteins, it matches or exceeds the 99.9th and even 99.99th percentiles of the DOCKSTRING database, including surpassing the best molecule present in the database itself ($N=260155$). Notably, TextGrad consistently underperforms Feedback Descent across all targets; while the TextGrad paper evaluated on similar DOCKSTRING tasks for only $\sim$10 iterations, we run both methods for 1000 steps and find that TextGrad’s pointwise conditioning does not scale well to high iteration budgets, confirming that accumulated feedback history is essential for sustained improvement. These findings show that Feedback Descent, a purely text-based method, can rival or outperform specialized graph-based algorithms, despite lacking handcrafted structural priors.
Figure˜5 shows optimization trajectories for PPARG. Feedback Descent achieves competitive trajectories relative to specialized methods, reaching high-scoring regions of chemical space with comparable or fewer oracle calls. This pattern holds across targets, suggesting that the method generalizes rather than relying on idiosyncrasies of a single protein system.
Analysis of discovered molecules. Figure˜6 illustrates the Pareto frontier between docking affinity (Vina score) and drug-likeness (QED) for PPARG. Feedback Descent recovers molecules that sit on or above the DOCKSTRING frontier, indicating that improvements in affinity are achieved without compromising drug-likeness. See Figure˜7 in the appendix for the full set of Pareto frontiers across all targets. These results show that feedback-guided search yields candidates that are not only potent but also balanced along multiple drug-relevant dimensions.
We also examine novelty by plotting Tversky similarity (CFP4 fingerprints) to approved DrugBank molecules against docking scores in Figure˜4. Across all targets, the correlations are weak or negative (Spearman $\rho$ between 0.39-0.39 and $0.40$), indicating that high-scoring candidates identified by Feedback Descent do not simply recycle functional groups from existing drugs, but instead explore novel regions of chemical space. For CDK2, no comparison is shown, as the target lacks any fully approved drugs in DrugBank with orthosteric binding as part of their mechanism of action, and thus does not satisfy our inclusion criteria.
Figure 5: Optimization trajectories on PPARG showing docking scores over oracle calls for Feedback Descent and specialized baselines. Feedback Descent quickly improves molecular docking scores within the first few hundred oracle calls.
Figure 6: Pareto frontier of docking affinity vs. drug-likeness, comparing Feedback Descent molecules (blue) to the DOCKSTRING database (gray). Feedback Descent finds novel molecules that meet or surpass known ones.
Feedback Descent Can Discover Novel Targeted Molecules Feedback Descent, operating in a purely textual form, consistently identifies novel molecules that surpass high-percentile baselines in DOCKSTRING. This demonstrates that iterative, feedback-guided optimization can enable models to genuinely explore unknown design spaces beyond their training distribution.
6 Discussion
This paper presents Feedback Descent, an inference-time framework that improves text artifacts through structured pairwise feedback. We validate it on visual design, prompt optimization, and molecule discovery, showing that text can serve as an optimizable medium, not just static data. Unlike parameter tuning, this approach can leverage richer textual signals, allowing for continual improvement without requiring retraining.
Limitations. The method relies on strong evaluators, which may be scarce in some domains. Training models to produce reliable feedback remains a prerequisite for harder tasks. For creative domains, strictly “following the gradient” may be limiting; balancing refinement with exploration is an important next step.
Acknowledgments
We would like to thank Wanqiao Xu, Allen Nie, Henrik Marklund, Mert Yuksekgonul, Anikait Singh, and Jubayer Ibn Hamid for helpful discussions and feedback. This work is supported by the OpenAI Superalignment Fellowship, the KFAS Fellowship, and the Schmidt Sciences AI2050 program.
Appendix A Formal Statements and Proofs
Proposition 1 (Linear convergence under PL with rationale-guided directions).
Let $r:Z\to\mathbb{R}$ be $L$-smooth and satisfy the $\mu$-PL condition (for maximization)
$\frac{1}{2}\|\nabla r(z)\|_{2}^{2}\;\geq\;\mu\big(r(z^{\star})-r(z)\big)\qquad\forall z\in Z.$
At iteration $t$, suppose a direction $v_{t}$ satisfies
$\mathbb{E}[\,v_{t}\mid z_{t}\,]\;=\;\alpha\,\nabla r(z_{t}),\qquad\mathbb{E}\!\left[\,\|v_{t}-\mathbb{E}[v_{t}\mid z_{t}]\|_{2}^{2}\,\big|\,z_{t}\right]\;\leq\;\sigma^{2}\|\nabla r(z_{t})\|_{2}^{2},$
with constants $\alpha>0$ and $\sigma\geq 0$, and define $\kappa_{1}\triangleq\alpha^{2}+\sigma^{2}$. Consider the update $z_{t+1}=z_{t}+\eta v_{t}$. If a constraint set $Z$ is present, assume $z_{t}+\eta v_{t}\in Z$ (i.e., the projection is inactive). With stepsize $\eta=\alpha/(L\kappa_{1})$,
$\mathbb{E}\!\left[r(z^{\star})-r(z_{t+1})\,\big|\,z_{t}\right]\;\leq\;\Big(1-\tfrac{\mu\,\alpha^{2}}{L\,\kappa_{1}}\Big)\,\big[r(z^{\star})-r(z_{t})\big].$
Unrolling yields
$\mathbb{E}\!\left[r(z^{\star})-r(z_{T})\right]\;\leq\;\Big(1-\tfrac{\mu\,\alpha^{2}}{L\,\kappa_{1}}\Big)^{T}\big[r(z^{\star})-r(z_{0})\big],$
so $\epsilon$-accuracy is achieved in
$T\;=\;O\!\left(\frac{L(\alpha^{2}+\sigma^{2})}{\mu\,\alpha^{2}}\,\log\frac{1}{\epsilon}\right)$
iterations.
Proof.
$L$-smoothness gives the two-sided bound
$r(z_{t}+\eta v_{t})\;\geq\;r(z_{t})+\eta\langle\nabla r(z_{t}),v_{t}\rangle-\tfrac{L}{2}\eta^{2}\|v_{t}\|_{2}^{2}.$
Taking conditional expectation and using $\mathbb{E}[v_{t}|z_{t}]=\alpha\nabla r(z_{t})$ and $\mathbb{E}\!\left[\|v_{t}\|_{2}^{2}\,\big|\,z_{t}\right]\leq(\alpha^{2}+\sigma^{2})\,\|\nabla r(z_{t})\|_{2}^{2}=\kappa_{1}\|\nabla r(z_{t})\|_{2}^{2}$,
$\mathbb{E}\!\left[r(z_{t+1})\mid z_{t}\right]\;\geq\;r(z_{t})+\Big(\eta\alpha-\tfrac{L}{2}\eta^{2}\kappa_{1}\Big)\,\|\nabla r(z_{t})\|_{2}^{2}.$
By the PL inequality, $\|\nabla r(z_{t})\|_{2}^{2}\geq 2\mu\,[r(z^{\star})-r(z_{t})]$, so
$\mathbb{E}\!\left[r(z^{\star})-r(z_{t+1})\mid z_{t}\right]\;\leq\;\Big(1-2\mu\eta\alpha+\mu L\eta^{2}\kappa_{1}\Big)\,[r(z^{\star})-r(z_{t})].$
Choosing $\eta=\alpha/(L\kappa_{1})$ makes the bracket equal to $1-\mu\alpha^{2}/(L\kappa_{1})$, yielding the claim. ∎
A.1 Query Complexity and Dimension Dependence
Dimension-Free Case. When rationales provide full gradient information ($v_{t}\in\mathbb{R}^{d}$) at unit cost, the query complexity equals $T$ and is dimension-independent:
$\text{Queries}=O\left(\frac{L(\alpha^{2}+\sigma^{2})}{\alpha^{2}\mu}\log\frac{1}{\epsilon}\right)$ (3)
Coordinate-Sparse Case. Suppose each query reveals one coordinate of $\nabla r(z_{t})$ chosen uniformly at random. Using the unbiased estimator $v_{t}=d\,(\partial_{i}r(z_{t}))\,e_{i}$ with $i\sim\mathrm{Unif}([d])$ gives $\alpha=1$, $\sigma^{2}=d-1$, and hence $\kappa_{1}=d$ and stepsize $\eta=1/(Ld)$. We have
$T=O\!\Big(\frac{Ld}{\mu}\log\frac{1}{\epsilon}\Big),\qquad\text{Queries}=O\!\Big(\frac{Ld}{\mu}\log\frac{1}{\epsilon}\Big).$
Equivalently, averaging $m$ independent coordinate queries per iteration yields $\sigma^{2}=(d-1)/m$; taking $m=d$ recovers $T=O((L/\mu)\log(1/\epsilon))$ with $d$ queries per iteration, so total queries remain $\Theta\!\big(\frac{Ld}{\mu}\log\frac{1}{\epsilon}\big)$.
This clarifies when and why dimension appears in the complexity.
Appendix B Lower Bounds for Exhaustive/Random Zeroth-Order Search
We formalize the intrinsic slowness of exhaustive (grid) search and best-of-$N$ random sampling when only function values (or preferences) are used without directional information. The hard instance is the strongly concave quadratic
$r(z)\;=\;r(z^{\star})-\tfrac{\mu}{2}\,\|z-z^{\star}\|_{2}^{2},\qquad z\in B_{R}(z^{\star})\subset\mathbb{R}^{d},$
whose $\epsilon$-optimal set is the ball $B_{\rho_{\epsilon}}(z^{\star})$ with radius $\rho_{\epsilon}=\sqrt{2\epsilon/\mu}$.
Proposition 2 (Grid-search lower bound).
Let $B_{R}(z^{\star})\subset\mathbb{R}^{d}$ and a hypercubic grid of spacing $h$. Its covering radius is $\rho=\frac{\sqrt{d}\,h}{2}$. To guarantee that for all placements of $z^{\star}$ there exists a grid point in the $\epsilon$-optimal ball $B_{\rho_{\epsilon}}(z^{\star})$ with $\rho_{\epsilon}=\sqrt{2\epsilon/\mu}$, it suffices that $\rho\leq\rho_{\epsilon}$ (i.e., $h\leq 2\rho_{\epsilon}/\sqrt{d}$). Furthermore, any such grid restricted to $B_{R}(z^{\star})$ must contain at least
$N\;\geq\;\left(\frac{R}{\rho}\right)^{\!d}\;=\;\Big(\frac{R\sqrt{d}}{2\rho_{\epsilon}}\Big)^{\!d}\;=\;\Big(\frac{\mu R^{2}d}{8\,\epsilon}\Big)^{\!d/2}$
points. Hence exhaustive grid search is exponential in $d$ and polynomial in $1/\epsilon$ with exponent $d/2$ on this family.
Proof.
Coverage of $B_{R}(z^{\star})$ by $N$ balls of radius $\rho$ centered at grid points implies $NV_{d}\rho^{d}\geq V_{d}R^{d}$, hence $N\geq(R/\rho)^{d}$. With $\rho=\sqrt{d}\,h/2$ and $h\leq 2\rho_{\epsilon}/\sqrt{d}$, we obtain $N\geq(R\sqrt{d}/(2\rho_{\epsilon}))^{d}$. Substitute $\rho_{\epsilon}=\sqrt{2\epsilon/\mu}$ to conclude. ∎
Proposition 3 (Best-of-$N$ random sampling lower bound).
Draw $X_{1},\dots,X_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}(B_{R}(z^{\star}))$ and let $\hat{z}=\arg\max_{i}r(X_{i})$ for $r(z)=r(z^{\star})-\tfrac{\mu}{2}\|z-z^{\star}\|_{2}^{2}$. Then with $a\triangleq 2/d$,
$\mathbb{E}\!\left[r(z^{\star})-r(\hat{z})\right]=\frac{\mu R^{2}}{2}\,N\,\mathrm{B}(1{+}a,N)=\frac{\mu R^{2}}{2}\,\Gamma(1{+}a)\,\frac{\Gamma(N+1)}{\Gamma(N+1+a)}.$
Moreover, for all $d\geq 1$ (so $a\in(0,2]$),
$\frac{\Gamma(N+1)}{\Gamma(N+1+a)}\ \geq\ (N+2)^{-a},$
and thus
$\mathbb{E}\!\left[r(z^{\star})-r(\hat{z})\right]\ \geq\ \frac{\mu R^{2}}{2}\,\Gamma\!\Big(1+\frac{2}{d}\Big)\,(N+2)^{-\frac{2}{d}}\ =\ \Omega\!\big(N^{-\frac{2}{d}}\big).$
Proof.
Let $R_{i}=\|X_{i}-z^{\star}\|_{2}$ and $R_{\min}=\min_{i}R_{i}$. The CDF of $R_{\min}$ is $F(r)=1-(1-(r/R)^{d})^{N}$ for $r\in[0,R]$. Differentiating, $f(r)=Ndr^{d-1}R^{-d}(1-(r/R)^{d})^{N-1}$. Then
$\mathbb{E}[R_{\min}^{2}]=\int_{0}^{R}r^{2}f(r)\,dr=NR^{2}\int_{0}^{1}t^{\frac{2}{d}}(1-t)^{N-1}dt=NR^{2}\,\mathrm{B}\!\big(1+\tfrac{2}{d},\,N\big),$
where $t=(r/R)^{d}$ and $\mathrm{B}$ is the Beta function. Using $\mathrm{B}(a,b)=\frac{\Gamma(a)\Gamma(b)}{\Gamma(a+b)}$ gives the exact expression. For the bound, we use the inequality $\Gamma(N+1)/\Gamma(N+1+a)\geq(N+2)^{-a}$ which holds for all $a\in(0,2]$ and $N\geq 1$. ∎
Appendix C Extended Experiment Section
C.1 Implementation Details
SVG Code Optimization.
We employ a tournament-style approach where gpt-5-mini generates SVG/TikZ code that gets rendered to PNG images for pairwise aesthetic comparisons by a separate instance of the same model acting as judge. The system maintains a “champion” design that only updates when both A-vs-B and B-vs-A orderings consistently agree on a winner, accumulating winning rationales into the generation prompt to guide aesthetic improvements across iterations. The judge provides natural language rationales explaining aesthetic preferences that inform subsequent generations.
IFBench Prompt Optimization.
We closely follow the setting of agrawal2025gepa for this experiment, including their choice of LLMs (Qwen3-8B and GPT-4.1-mini) and multi-stage DSPy programs. For Qwen3-8B, we use 10 iterations with no early stopping (patience=0) with temperature 0.6, and evaluate 2 prompt candidates per round across 20 hard examples. For GPT-4.1-mini, we increase to 15 iterations with early stopping (patience=5) with temperature 1.0 for exploration. Both configurations use the same prompt improver template, described in Section˜C.3.
We programmatically generate a textual description of the difference between two prompts. To compare two prompts, we first partition the training set into four quadrants based on outcomes: examples where prompt A succeeds and B fails (A_wins), A fails and B succeeds (B_wins), both fail (tie_fail), or both succeed (tie_success). We then use the same LLM to propose hypotheses about input characteristics (based on the prompt text) and output patterns (based on the response and evaluation feedback), producing around 20 hypotheses for each category. To evaluate whether each hypothesis applies to each example, we use the same LLM as the tagger that outputs binary labels (1 if the hypothesis matches, 0 otherwise) for all hypotheses in a single call per example, processing hundreds of examples in parallel. We then compute lift metrics for each hypothesis-quadrant pair, where lift is the ratio of conditional probability to the base rate (i.e., how much more likely an outcome is given the hypothesis holds). We validate the hypotheses statistically using Fisher’s exact test, and filter for hypotheses that are statistically significant at $p<0.1$ with minimum support of 3 examples and lift $\geq 2.0$ for A/B wins or $\geq 1.5$ for failures. This analysis identifies which input patterns correlate with differential performance and which output characteristics appear when one prompt outperforms the other, providing actionable insights for prompt improvement.
Molecule Optimization.
We implement molecular optimization using the DOCKSTRING package [garciaortegon2021dockstring] for protein-ligand docking simulations across six therapeutic targets. The system begins with three simple seed molecules (acetamide, pentane, benzene) and progressively evolves SMILES strings through iterative feedback loops that incorporate RDKit molecular properties, protein binding site information, and similarity comparisons to approved drugs as metadata. We use the combined score function suggested by DOCKSTRING:
$s_{\text{overall}}(\text{molecule},\text{protein})=-\texttt{Vina}(\text{molecule},\text{protein})-10*(1-\texttt{QED}(\text{molecule})),$ (4)
where Vina provides the binding affinity prediction (kcal/mol, more negative is better) and the QED penalty term penalizes molecules with poor drug-likeness, with lower overall scores indicating better molecules that balance binding strength and drug-like properties. Note that QED scores range from 0 to $1$ while Vina scores typically range from 3.0-3.0 to 12.0-12.0 kcal/mol. For Feedback Descent, we use a batch size of 8 and top-k selection of 10 examples.
C.2 Additional Results
Figure 7: Pareto frontiers of discovered molecules (blue) compared against molecules in the DOCKSTRING dataset (gray) across six protein targets. The highlighted orange markers indicate molecules on the discovered Pareto frontier, achieving joint improvements in docking affinity (Vina score) and drug-likeness (QED).
Figure˜7 shows that across all protein targets, the discovered molecules extend beyond the DOCKSTRING baseline along both axes. The resulting Pareto frontiers illustrate consistent improvements in the joint trade-off between docking affinity and drug-likeness, highlighting that feedback-guided search yields coordinated gains rather than isolated outliers.
C.3 Prompt Templates
We use the following prompt for the judge for the Anatomy SVG task. The rubrics for the other tasks are written in a similar style, translating a particular aesthetic into operational rules that minimize ambiguity.
Anatomy Judge Rubric We use the following prompt templates for candidate generation and rationale generation for prompt optimization. System Prompt Template for Prompt Optimization System Prompt Template for Molecule Optimization Example of Protein Metadata (ADRB1) Example of Molecule Metadata (CCCCC)